Sweetwater Company manufactures two products, Mountain Mist and Valley Stream. The company prepares its master budget on the basis of standard costs The following data are for March: Mountain Mist Valley Stream Direct materials Direct labor Variable overhead (per direct labor-hour) Fixed overhead (per month) Expected activity (direct labor-hours) 3 ounces at $14.60 per ounce 5 hours at $60.40 per hour 548 $347 475 6.150 4 ounces at $16.90 per ounoe 6 hours at $79 per hour $52.90 $398,580 7,800 Actual results Direct material (purchased and used) Direct labor Variable overhead Fixed overhead Units produced (actual) 3.500 ounces at $13.90 per ounce 4,940 hours at $61.75 per hour $254550 $317.950 040 units 4,800 ounces at $18.25 per ounce 7.450 hours at $79.60 per hour 382,510 5396,100 1,240 units Required a. Compute a variance analysis for each variable cost for each product. (Do not round intermediate calculations. Indicate the effect of each variance by selecting F for favorable, or U for unfavorable. If there is no effect, do not select either option.) Price Price Direct materials Direct labor Variable overhead b. Compute a fixed overhead variance analysis for each product. (Do not round intermediate calculations. Indicate the effect of each variance by selecting F for favorable, or U for unfavorable. If there is no effect, do not select either option.) Stream Price Variance Price Variance Fixed overhead
Answer #1
a) Mountain Mist Valley stream
Price Variance Efficiency Variance Price Variance Efficiency Variance
Direct Materials 2450 F 5548 U 6480 U 2704 F
=-3500*(13.9-14.6) =(3500-1040*3)*14.6 =4800*(18.25-16.9) =-(4800-1240*4)*16.9
Direct labor 6669 U 15704 F 4470 U 790 U
=4940*(61.75-60.4) =-(4940-1040*5)*60.4 =7450*(79.6-79) =(7450-1240*6)*79
Variable overhead 17430 U 5280 F 11595 F 790 U
=254550-4940*48 =-(4940-1010*5)*48 =-(382510-7450*52.9) =(7450-1240*6)*79
b) Mountain Mist Valley stream
Price Variance Production Volume Variance Price Variance Production Volume Variance
Fixed overhead 29525 F 53675 U 2480 F 18396 U
=-(317950-347475) =347475-1040*5*347475/6150 =-396100+398580 =398580-1240*6*398580/7800